OpenStudy (anonymous):

DO the following: A. solve the equation B. answer the question Deni can clean the pool four times as fast as her sister lamur. Together they can do the job in 3 hours. How long would it take each of them working alone?

OpenStudy (anonymous):

D = Deni L = Lamur h = hours Dh = 4Lh D + L = 3h

OpenStudy (anonymous):

Cool I got that :)

OpenStudy (anonymous):

So substitute... 4Lh + L = 3h so that L = 3/5

OpenStudy (anonymous):

Now solve for D by substituting 3/5 in D = 4L

OpenStudy (anonymous):

D = 3/5(4)

OpenStudy (anonymous):

d= 2.4

OpenStudy (anonymous):

Then switch the D and L for the answer.

OpenStudy (anonymous):

so do 4(3/5) ?

OpenStudy (anonymous):

So L = 2.4h and D = 3/5h